Solve each of the following equations for x.

2/5 = x-2/20

Answer by checkley71(8403)

I'm not sure if I explained it correctly, 2/5 = x-2 divided by 20.
I'LL TRY THIS ONE
2/5=(X-2)/20 YOU NEED TO MAKE MORE USE OF THE ().
CROSS MULTIPLY
2*20=5(X-2)
40=5X-10
5X=40-10
5X=30
X=6 ANSWER.
-------------------------------------------------------------
2/5=X-2/20
X=2/5+2/20
X=(2*4+2)/20
X=(8+2)/20
X=10/20
X=1/2 ANSWER.
PROOF
.4=.5-.1
.4=.4